John and Mary are skating at an ice rink. John skates at a constant speed of 6.6 m/s, with respect to the ice surface, directly south. Mary skates at a constant speed of 12.3 m/s in a direction of 33° west of south, with respect to the ice. What are the magnitude and direction of Mary's velocity with respect to John? Give the direction as measured relative to due south.

What is magnitude

What is direction
Answer #1

West component of Mary's speed = (sin 33 x 12.3) = 6.7 m/sec.
South component = (cos33 x 12.3) = 10.3 m/sec.
South component = (10.3 - 6.6) = 3.7 m/sec. relative to John.
The west component is 6.7 m/sec. relative to John.
Relative V = sqrt.(6.7^2 + 3.7^2) = 7.65 m/sec.
Direction perceived by John = arctan (6.7/3.7) = 61.1 degrees west of his direction south.